OpenStudy (anonymous):

Please Help!! Suppose you have a box with 3 blue marbles, 2 red marbles, and 4 yellow marbles. You are going to pull out one marble, record its color, put it back in the box and draw another marble. What is the probability of pulling out a red marble followed by a blue marble? Be sure to reduce your answers. a. 6/81 = 2/27 b. 6/18 = 1/3 c. 5/9 d. 5/18

OpenStudy (misssunshinexxoxo):

Which do you believe it could be?

OpenStudy (anonymous):

Give me a second I have to find the answer I picked

OpenStudy (misssunshinexxoxo):

These probabilities in the beginning could get tricky

OpenStudy (anonymous):

Hmm I can't find it

OpenStudy (misssunshinexxoxo):

P(red) = 2/9 P(blue) = 3/9

OpenStudy (misssunshinexxoxo):

P(redblue) = (2/9)(3/9) = 6/81 = 2/27

OpenStudy (decentnabeel):

The multiplication rule says we need to find P(red) P(blue). P(red) = 2/9 P(blue) = 3/9 P(redblue) = (2/9)(3/9) = 6/81 = 2/27

OpenStudy (anonymous):

Well, you first have to find the chance of getting a red marble, then multiply it by the chance of blue. 2/9*3/9=6/81=2/27

OpenStudy (misssunshinexxoxo):

Well done all. :)